### v1

Abstract
In sensitive application domains, multi-objective hyperparameter selection can ensure the reliability of AI models prior to deployment, while optimizing auxiliary performance metrics. The state-of-the-art Pareto Testing (PT) method guarantees statistical reliability constraints by adopting a multiple hypothesis testing framework. In PT, hyperparameters are validated one at a time, following a data-driven order determined by expected reliability levels. This paper introduces a novel framework for multi-objective hyperparameter selection that captures the interdependencies among the reliability levels of different hyperparameter configurations using a directed acyclic graph (DAG), which is termed the reliability graph (RG). The RG is constructed based on prior information and data by using the Bradley-Terry model. The proposed approach, RG-based PT (RG-PT), leverages the RG to enable the efficient, parallel testing of multiple hyperparameters at the same reliability level. By integrating False Discovery Rate (FDR) control, RG-PT ensures robust statistical reliability guarantees and is shown via experiments across diverse domains to consistently yield superior solutions for multi-objective calibration problems.
1 Introduction
1.1 Context and Motivation
Hyperparameter optimization is an essential element of the machine learning pipeline, particularly in sensitive domains such as healthcare, finance, and engineering Bischl et al. (2023); Yang & Shami (2020). Hyperparameters to be set prior to deployment include training parameters such as learning rates, regularization constants, and reward priorities in reinforcement learning systems Eimer et al. (2023), inference parameters, such as decision thresholds Hancock et al. (2022), and complexity settings determining numerical accuracy and computational efficiency Hu et al. (2021). Hyperparameter selection typically targets multiple performance objectives, such as classification accuracy, calibration, fairness, and domain-specific criteria, such as tracking precision Dong et al. (2018), perplexity Rahnama et al. (2018), and transmission rates Chen et al. (2020).
Figure 1: Comparison of Pareto Testing (PT) Laufer-Goldshtein et al. (2023) and reliability graph-based PT (RG-PT) proposed in this work. (a) PT first estimates the Pareto front in the space of risk functions to be controlled (horizontal axis) and optimized (vertical axis), with the red dashed line indicating the reliability constraint. (b) PT then tests the reliability of hyperparameters $\lambda$ sequentially, in the order of estimated reliability. (c) RG-PT first estimates interdependencies among the reliability levels of different hyperparameter configurations using a directed acyclic graph (DAG), which is termed the reliability graph (RG). This enables efficient parallel testing of multiple hyperparameters, potentially discovering more effective hyperparameter configurations (see Fig. 2).
Validation can be costly. Notably, determining training or fine-tuning parameters in large language models (LLMs) requires running expensive optimization procedures, and many engineering problems require interaction with the real world Polese et al. (2023). Thus, in modern settings, it is common to limit the search to a pre-selected discrete set of configurations for the hyperparameters Bergstra & Bengio (2012).
Most existing hyperparameter selection strategies make a best-effort attempt at addressing multi-objective problems. This is done by estimating functions of the Pareto front for the performance criteria of interest Deb et al. (2002); Hernández-Lobato et al. (2016); Paria et al. (2020); Li et al. (2020). An important, recent, exception is Pareto Testing (PT) Laufer-Goldshtein et al. (2023). PT builds on Learn-Then-Test (LTT) Angelopoulos et al. (2021), which pioneered the use of multiple hypothesis testing (MHT) to ensure the statistical validity of hyperparameter selection.
As illustrated in Fig. 1, PT carries out a preliminary best-effort optimization step to estimate the Pareto front in the space of the objectives of interest. These encompass risk functions that must be controlled, encoding reliability constraint, as well as auxiliary risk functions to be optimized. PT proceeds by testing the candidate hyperparameters according to the linear order determined by the increasing value of the estimated reliability risk functions (see Fig. 1(b)).
This paper is motivated by the observation that, especially in engineering applications, the hyperparameter space is often highly structured, and not all configurations can be neatly ordered by their reliability levels. To account for this, we introduce a novel multi-objective hyperparameter selection method that operates on a directed acyclic graph (DAG), termed the reliability graph (RG). In an RG, nodes represent candidate hyperparameter configurations and an edge directed from one hyperparameter to another indicates that the reliability of the first hyperparameter is predictive of the reliability level of the second hyperparameter (see Fig. 1(c)).
The proposed hyperparameter selection method, referred to as RG-based Pareto Testing (RG-PT), constructs an RG based on prior knowledge and limited validation data. Prior knowledge may include properties such as the reduction in reliability that may result from an increased sparsity at the levels of model parameters Zhou et al. (2021) or activations Chen et al. (2023). By incorporating MHT mechanisms operating on a DAG Ramdas et al. (2019) within the PT framework, RG-PT allows for a more efficient testing that can discover more effective hyperparameter configurations, while offering reliability guarantees (see Fig. 2).
1.2 Further Related Work
Hyperparameter Selection: Traditional methods for hyperparameter optimization like grid search and random search Bergstra & Bengio (2012) are widely used but they can be inefficient. In contrast, more advanced techniques such as Bayesian optimization Snoek et al. (2012) and bandit-based methods, such as Hyperband Li et al. (2018) and gradient-based optimization Maclaurin et al. (2015), can enhance computational efficiency via resource allocation, but they lack statistical guarantees. This gap addressed is by the mentioned frameworks incorporating MHT Angelopoulos et al. (2021); Laufer-Goldshtein et al. (2023).
Multi-Objective Optimization: Modern AI applications often require optimizing multiple objectives such as accuracy and efficiency. This can be formally done through Pareto optimization to identify all the feasible trade-off points among different objectives Deb et al. (2002). PT Laufer-Goldshtein et al. (2023) applies MHT to provide statistical reliability while balancing multiple objectives.
Conformal Prediction and LTT: Conformal prediction constructs statistically valid prediction intervals using calibration data Shafer & Vovk (2008); Romano et al. (2019). Drawing inspiration from conformal prediction, LTT Angelopoulos et al. (2021) frames hyperparameter selection as an MHT problem. While Angelopoulos et al. (2021) mentions the possible use of graph-based approaches, these are limited to fixed user-defined graphs or linear directed graphs (chains). Other extensions of LTT include quantile control Farzaneh et al. (2024), applications to information theory Farzaneh & Simeone (2024), and adaptive testing mechanisms Zecchin & Simeone (2024).
1.3 Main Contributions
Figure 2: Illustration of the benefits of the proposed RG-PT hyperparameter selection scheme over the state-of-the-art LTT and PT for an object detection application Angelopoulos et al. (2021). The red arrows mark the objects not detected by an object recognition model calibrated using LTT or PT that are instead detected by the same model calibrated via RG-PT (see Sec. 5.2 for details).
The main contributions of this paper are as follows.
Methodology: We propose RG-PT, a novel multi-objective hyperparameter selection framework that systematically utilizes interdependencies among the reliability levels of candidate hyperparameter configurations. RG-PT infers precedence relationships among the reliability levels of hyperparameters from prior information and data by constructing an RG via the Bradley-Terry (BT) ranking model Hunter (2004). Then, it applies MHT-based hyperparameter selection on the RG by following the graphical testing method in Ramdas et al. (2019).
Applications: We demonstrate the effectiveness of RG-PT through experiments in various scenarios, including language model calibration Peters & Martins (2021), object detection Angelopoulos et al. (2021), image classification Franceschi et al. (2024), and telecommunications engineering Valcarce (2020), highlighting its advantages over conventional methods.
The rest of this paper is organized as follows. In Sec. 2, we define the multi-objective hyperparameter selection problem. Sec. 3 provides an overview of PT, and Sec. 4 details the proposed RG-PT framework. Experimental results are presented in Sec. 5. We conclude the paper in Sec. 6.
2 Multi-Objective Hyperparameter Selection
In this section, we define the problem of multi-objective hyperparameter selection, and we show how this problem can be formulated via MHT by following references Angelopoulos et al. (2021); Laufer-Goldshtein et al. (2023).
2.1 Problem Definition
Consider a predefined discrete and finite set $\Lambda$ of hyperparameters $\lambda$, which govern the performance of a machine learning model (see, e.g., Bergstra & Bengio (2012)). Hyperparameter selection aims at obtaining a hyperparameter $\lambda\in\Lambda$ by considering its performance via $L$ risk functions. Specifically, when tested on a data point $Z$, a hyperparameter $\lambda$ attains risk values $r_{l}(Z,\lambda)$ for $l=1,\ldots,L$. The risk functions $r_{l}(Z,\lambda)$ are negatively oriented, meaning that lower risk values correspond to better-performing hyperparameters. The risks are normalized within the range $0\leq r_{l}(Z,\lambda)\leq 1$.
For each performance criterion $l=1,\ldots,L$, the average risk function is defined as
$R_{l}(\lambda)=\mathbb{E}_{Z}\left[r_{l}(Z,\lambda)\right],$ (1)
where the expectation is taken over the distribution $P_{Z}$ of the data $Z$.
We partition the set of $L$ risk functions into the following two groups:
1. Reliability risk functions: The first set of risk functions $\{R_{l}(\lambda)\}_{l=1}^{L_{c}}$ must be controlled via the choice of the hyperparameter $\lambda$. In particular, a hyperparameter configuration $\lambda$ is said to be reliable if it guarantees the constraints
$R_{l}(\lambda)\leq\alpha_{l}\quad\text{for all}\quad l=1,\ldots,L_{c}.$ (2)
2. Auxiliary risk functions: The second set of performance measures $\{R_{l}(\lambda)\}_{l=L_{c}+1}^{L}$ are unconstrained, and they are ideally optimized via the selection of the hyperparameter $\lambda$.
Accordingly, the goal of hyperparameter selection is defined as the multi-objective problem
$\min_{\lambda\in\Lambda}\;\{R_{L_{c}+1}(\lambda),R_{L_{c}+2}(\lambda),\dots,R_{L}(\lambda)\}$ (3)
subject $\text{subject to}\;R_{l}(\lambda)<\alpha_{l}\;\text{for all}\;1\leq l\leq L_{c},$
which targets the minimization of the auxiliary risk functions $\{R_{l}(\lambda)\}_{l=L_{c}+1}^{L}$ under constraints on the reliability risk functions $\{R_{l}(\lambda)\}_{l=1}^{L_{c}}$. Solving a multi-objective optimization problem such as (3) may entail identifying the entire Pareto front of dominant solutions $\lambda\in\Lambda$, or obtaining specific solutions corresponding to scalar criteria Deb et al. (2016); Jamieson & Talwalkar (2016).
The problem (3) cannot be directly addressed since the data distribution $P_{Z}$ is unknown. However, we assume to have access to i.i.d. data $\mathcal{Z}=\{Z_{j}\}_{j=1}^{n}$ drawn from the unknown data distribution $P_{Z}$. For any subset $\widetilde{\mathcal{Z}}\subseteq\mathcal{Z}$, we write as
$\hat{R}_{l}(\lambda|\widetilde{\mathcal{Z}})=\frac{1}{|\widetilde{\mathcal{Z}}|}\sum_{Z\in\widetilde{\mathcal{Z}}}r_{l}(Z,\lambda)$ (4)
the corresponding empirical estimate of risk function $R_{l}(\lambda)$.
2.2 Hyperparameter Selection as Multiple Hypothesis Testing
As proposed in Angelopoulos et al. (2021), hyperparameter selection can be formally addressed as an MHT problem. Accordingly, for each hyperparameter $\lambda\in\Lambda$, we define the null hypothesis $\mathcal{H}_{\lambda}$ that hyperparameter $\lambda$ is not reliable, violating the constraints (2), i.e.,
$\mathcal{H}_{\lambda}:\;\text{there exists}\;l\in\{1,\dots,L_{c}\}\text{ such that }R_{l}(\lambda)>\alpha_{l}.$ (5)
Thus, rejecting the null hypothesis $\mathcal{H}_{\lambda}$ implies that hyperparameter $\lambda$ is reliable, meeting all the constraints in (2). A rejection is also referred to as a discovery. A discovery is false if the selected hyperparameter $\lambda$ is actually unreliable, satisfying the null hypothesis $\mathcal{H}_{\lambda}$.
As discussed in the next section, by formulating hyperparameter selection as an MHT problem, we can leverage statistical tools that guarantee false discovery rate (FDR) requirements Casella & Berger (2024), ensuring that any selected configuration satisfies the risk constraints in (3) with a prescribed probability.
3 Background: Pareto Testing
In this section, we review PT Laufer-Goldshtein et al. (2023), a state-of-the-art MHT-based multi-objective hyperparameter selection method.
3.1 Statistical Reliability Constraints
Problem (3) cannot be directly solved since we only assume access to the empirical estimates. Following LTT Angelopoulos et al. (2021), using the data set $\mathcal{Z}$, PT finds a subset $\hat{\Lambda}_{\mathcal{Z}}\subseteq\Lambda$ of hyperparameters that satisfies the statistical constraint
$\mathrm{Pr}_{\mathcal{Z}}\big{[}R_{l}(\lambda)\leq\alpha_{l}$ for all $\text{for all }l=1,\ldots,L_{c},$ (6)
$\text{and for all }\lambda\in\hat{\Lambda}_{\mathcal{Z}}\big{]}\geq 1-\delta,$
where $0\leq\delta\leq 1$ is a user-specified probability of failure, and the probability is evaluated over the calibration data $\mathcal{Z}$. The condition (6), known as family-wise error rate (FWER) control, ensures that all the hyperparameters $\lambda$ in the selected set $\hat{\Lambda}_{\mathcal{Z}}$ are reliable with probability at least $1-\delta$.
3.2 Pareto Testing
PT Laufer-Goldshtein et al. (2023) leverages held-out data to select the testing order for addressing problems of the form (3). The procedure is summarized in Fig. 3, and it consists of the following steps:
Estimating the Pareto front for all risk measures: PT divides the calibration dataset $\mathcal{Z}$ into two disjoint subsets, $\mathcal{Z}_{\text{OPT}}$ and $\mathcal{Z}_{\text{MHT}}$. Using the data set $\mathcal{Z}_{\text{OPT}}$, it identifies the subset $\Lambda_{\text{OPT}}\subseteq\Lambda$ of hyperparameters that are on the Pareto front of the space of estimated risk measures $\{\hat{R}_{l}(\lambda|\mathcal{Z}_{\text{OPT}})\}_{l=1}^{L}$. This is done by addressing the multi-objective optimization problem (3) with the estimates $\{\hat{R}_{l}(\lambda|\mathcal{Z}_{\text{OPT}})\}_{l=1}^{L}$ in lieu of the true risks $\{R_{l}(\lambda|\mathcal{Z}_{\text{OPT}})\}_{l=1}^{L}$ by using any existing algorithm Laufer-Goldshtein et al. (2023).
Ordering the hyperparameters: PT creates a global linear ordering on the hyperparameters in the set $\Lambda_{\text{OPT}}$ by using the estimates $\{\hat{R}_{l}(\lambda|\mathcal{Z}_{\text{OPT}})\}_{l=1}^{L_{c}}$ of the reliability risk functions.
FWER-controlling MHT: Using the data set $\mathcal{Z}_{\text{MHT}}$, FWER-controlling MHT is carried out via fixed sequence testing (FST) Maurer (1995) by following the order found in the previous step. This results in the subset $\hat{\Lambda}_{\mathcal{Z}}$ of reliable hyperparameters that are guaranteed to satisfy the condition (6).
Addressing the multi-objective optimization problem: Given the subset $\hat{\Lambda}_{\mathcal{Z}}$, PT addresses the problem
$\min_{\lambda\in\hat{\Lambda}_{\mathcal{Z}}}\{\hat{R}_{L_{c}+1}(\lambda|\mathcal{Z}_{\text{OPT}}),\ldots,\hat{R}_{L}(\lambda|\mathcal{Z}_{\text{OPT}})\},$ (7)
where the auxiliary risk functions $R_{l}(\lambda)$ in (3) are replaced with the corresponding empirical estimates (4) obtained with data set $\mathcal{Z}_{\text{OPT}}$.
To elaborate on steps ②-③, for each null hypothesis $\mathcal{H}_{\lambda}$ in (5), given a data set $\widetilde{\mathcal{Z}}\subseteq\mathcal{Z}$, define the $p$-value Rice (2006)
$p_{\lambda,l}(\widetilde{\mathcal{Z}})=e^{-2|\widetilde{\mathcal{Z}}|(\alpha_{l}-\hat{R}_{l}(\lambda|\widetilde{\mathcal{Z}}))_{+}^{2}}$ (8)
for $l=1,\ldots,L_{c}$, and the combined $p$-value
$p_{\lambda}(\widetilde{\mathcal{Z}})=\max_{1\leq l\leq L_{c}}p_{\lambda,l}(\widetilde{\mathcal{Z}}).$ (9)
In step ②, hyperparameters in set $\Lambda_{\text{OPT}}$ are ordered from low to high based on the combined $p$-value $p_{\lambda}(\mathcal{Z}_{\text{OPT}})$, reflecting an ordering from most to least reliable hyperparameters according to data $\mathcal{Z}_{\text{OPT}}$. FST, a sequential MHT algorithm (see Appendix A), is then applied by using the combined $p$-values $p_{\lambda}(\mathcal{Z}_{\text{MHT}})$.
Figure 3: Illustration of the operation of PT Laufer-Goldshtein et al. (2023) for two risk functions $R_{1}(\lambda)$ and $R_{2}(\lambda)$. The objective is to control the risk $R_{1}(\lambda)$ to be below the threshold $\alpha$, while minimizing the risk $R_{2}(\lambda)$: ① Approximate the Pareto frontier using the estimates $\hat{R}_{1}(\lambda|\mathcal{Z}_{\text{OPT}})$ and $\hat{R}_{2}(\lambda|\mathcal{Z}_{\text{OPT}})$ to form the subset of hyperparameters $\Lambda_{\text{OPT}}$; ②-④ Apply a sequential FDR-controlling algorithm on the hyperparameters in $\Lambda_{\text{OPT}}$ to obtain the subset $\hat{\Lambda}_{\mathcal{Z}}$, and select the hyperparameter $\lambda^{*}$ that minimizes $\hat{R}_{2}(\lambda|\mathcal{Z}_{\text{OPT}})$.
4 Reliability Graph-Based Pareto Testing
In this section, we introduce RG-PT, a novel hyperparameter selection strategy based on MHT that adopts a testing schedule based on the novel concept of RG.
4.1 Overview
The general principle underlying RG-PT is that the reliability of some hyperparameters can be highly predictive of the reliability of other hyperparameters. This structure is encoded in an RG by leveraging prior information and data via the BT model Hunter (2004). By testing on a DAG, RG-PT supports a more flexible and efficient testing procedure that allows for the testing of multiple hyperparameters in parallel.
Using a partition $\{\mathcal{Z}_{\text{OPT}},\mathcal{Z}_{\text{MHT}}\}$ of the data set $\mathcal{Z}$, at a high level, RG-PT consists of the following steps, which are illustrated in Fig. 4:
Estimating the Pareto front for all risk measures: Using the data set $\mathcal{Z}_{\text{OPT}}$, RG-PT identifies the the subset $\Lambda_{\text{OPT}}\subseteq\Lambda$ of hyperparameters that are estimated to lie on the Pareto front in the same manner as PT (see previous section).
Learning the reliability graph: Rather than ordering the hyperparameters globally and sequentially as done by PT, RG-PT creates an RG, with nodes given by the hyperparameters in subset $\Lambda_{\text{OPT}}$. This is done by following the principle that hyperparameters $\lambda\in\Lambda_{\text{OPT}}$ whose reliability levels are predictive of the reliability levels of other hyperparameters $\Lambda^{\prime}\subset\Lambda_{\text{OPT}}$ should be tested before the hyperparameters $\Lambda^{\prime}$. Details are provided in Sec. 4.2.
FDR-controlling MHT: FDR-controlling MHT is carried out by following the order encoded by the RG. As explained in Sec. 4.2, this is done by using DAGGER Ramdas et al. (2019), targeting FDR requirements.
Addressing the multi-objective optimization problem: Using the estimated subset $\hat{\Lambda}_{\mathcal{Z}}$ of reliable hyperparameters, the multi-objective optimization problem (7) is addressed in the same manner as PT (see previous section).
Figure 4: Illustration of the main steps of RG-PT: ① Estimate the hyperparameters $\Lambda_{\text{OPT}}$ lying on the Pareto front; ② Build the RG over the selected hyperparameters; ③ Apply an FDR-controlling MHT procedure, DAGGER, to the RG.
4.2 Learning the Reliability Graph
After obtaining the estimated Pareto front $\Lambda_{\text{OPT}}$, RG-PT constructs an RG to encode the expected relationships between the reliability levels attained by the candidate hyperparameters in the set $\Lambda_{\text{OPT}}$. The RG is a DAG, in which each node represents a hyperparameter, and edges are directed to describe a reliability hierarchy. Specifically, edges encode the expectation that parent nodes are predictive of the reliability of their child nodes.
Starting from the nodes with no parents and following the direction of the edges in the RG, one encounters hyperparameters that are estimated to be increasingly unreliable. Generalizing the linear ordering assumed by PT, the DAG structure adopted by RG-PT can assign the same expected reliability ranking to multiple hyperparameters. Specifically, all the hyperparameters at the same depth in the DAG are deemed to have the same relative reliability level.
In order to learn the RG, RG-PT leverages the data set $\mathcal{Z}_{\text{OPT}}$, as well as, possibly, prior information about the relative reliability of pairs of hyperparameters. This is done by following two steps:
1. Depth assignment: The hyperparameters in set $\Lambda_{\text{OPT}}$ are ranked in terms of their expected reliability, while allowing for multiple hyperparameters to be ranked equally. This effectively assigns a depth level $d$ in the DAG to each hyperparameter, with multiple hyperparameters possibly sharing the same depth level. This step is explained in Sec. 4.2.1.
2. Learning the directed edges: Given any hyperparameter $\lambda$ at some depth $d$, RG-PT selects a subset of hyperparameters at the previous depth level $d-1$ to serve as parents of the hyperparameter $\lambda$. This is done by choosing the hyperparameters at depth $d-1$ that are most predictive of the reliability level of hyperparameter $\lambda$. This step is explained in Sec. 4.2.2.
4.2.1 Depth Assignment
Fix a number $D\leq|\Lambda_{\text{OPT}}|$ of levels for the DAG. With $D=|\Lambda_{\text{OPT}}|$, one can assign each hyperparameter $\lambda\in\Lambda_{\text{OPT}}$ a distinct level, yielding a global ordering and recovering PT. Conversely, with $D=1$, all hyperparameters $\lambda\in\Lambda_{\text{OPT}}$ are assigned to the same level. The setting of interest is thus $1<D<|\Lambda_{\text{OPT}}|$, which is assumed from now on.
Depth assignment is carried out by first obtaining a score $s(\lambda)$ for all hyperparameters $\lambda\in\Lambda_{\text{OPT}}$ using the data set $\mathcal{Z}_{\text{OPT}}$, and then partitioning the set $\Lambda_{\text{OPT}}$ into $D$ clusters according to the obtained scores.
To compute the scores $s(\lambda)$ for hyperparameters $\lambda\in\Lambda_{\text{OPT}}$, we use the BT model Hunter (2004). The BT model converts pairwise counts $w_{ij}$ for all pairs of hyperparameters $\lambda_{i}$ and $\lambda_{j}$ in subset $\Lambda_{\text{OPT}}$ into per-hyperparameter scores $s(\lambda)$ for all $\lambda\in\Lambda_{\text{OPT}}$. The pairwise counts measure the number of times that hyperparameter $\lambda_{i}$ was found to be more reliable than hyperparameter $\lambda_{j}$. In RG-PT, we propose to evaluate the pairwise counts $w_{ij}$ by leveraging two sources of information:
$\bullet$ Prior information: Prior information is encoded by pairwise probabilities $0\leq\eta_{ij}\leq 1$ for each pair of hyperparameters $\lambda_{i},\lambda_{j}\in\Lambda_{\text{OPT}}$. This probability reflects the expected rate at which hyperparameter $\lambda_{i}$ is observed to be more reliable than hyperparameter $\lambda_{j}$. Note that we have $\eta_{ji}=1-\eta_{ij}$. The strength of the prior information is determined via a pseudocount variable $n_{p}$ as in the standard categorical-Dirichlet model Bishop & Nasrabadi (2006). A larger pseudocount $n_{p}$ indicates a stronger trust in the prior information. Importantly, the statistical guarantees of RG-PT do not depend on the choice of the prior probabilities $\{\eta_{ij}\}$ and pseudocount $n_{p}$, which can, however, improve the capacity of RG-PT to optimize the auxiliary risk functions. Note that in the absence of prior information, one can set $n_{p}=0$.
$\bullet$ Data: Using the $p$-values $p_{\lambda_{i}}(\mathcal{Z}_{\text{OPT}})$ and $p_{\lambda_{j}}(\mathcal{Z}_{\text{OPT}})$ in (9), we evaluate the data-driven probability
$p_{ij}(\mathcal{Z}_{\text{OPT}})=\frac{p_{\lambda_{i}}(\mathcal{Z}_{\text{OPT}})}{p_{\lambda_{i}}(\mathcal{Z}_{\text{OPT}})+p_{\lambda_{j}}(\mathcal{Z}_{\text{OPT}})}$ (10)
that hyperparameter $\lambda_{i}$ is more reliable than hyperparameter $\lambda_{j}$. Note that we have $p_{ij}(\mathcal{Z}_{\text{OPT}})=1-p_{ji}(\mathcal{Z}_{\text{OPT}})$.
Overall, the pair-wise probability $w_{ij}$ is obtained by combining prior information and data as
$w_{ij}=|\mathcal{Z}_{\text{OPT}}|p_{ij}(\mathcal{Z}_{\text{OPT}})+n_{p}\eta_{ij},$ (11)
so that the strength of the prior depends on the ratio $n_{p}/|\mathcal{Z}_{\text{OPT}}|$ between the pseudocount $n_{p}$ and the number of data points $|\mathcal{Z}_{\text{OPT}}|$.
Using the BT model, the scores $s(\lambda_{i})$ for all hyperparameters $\lambda_{i}\in\Lambda_{\text{OPT}}$ are obtained by maximizing the log-likelihood Hunter (2004)
$\sum_{i=1}^{|\Lambda_{\text{OPT}}|}\sum_{j=1}^{|\Lambda_{\text{OPT}}|}\left(w_{ij}\ln\left(\frac{s(\lambda_{i})}{s(\lambda_{i})+s(\lambda_{j})}\right)\right),$ (12)
with $w_{ii}=0$ for all $1\leq i\leq|\Lambda_{\text{OPT}}|$. With this design, in the absence of prior information ($n_{p}=0$), the BT model reduces to assigning scores directly proportional to the $p$-values $p_{\lambda}(\mathcal{Z}_{\text{OPT}})$, as is done in PT Laufer-Goldshtein et al. (2023). Therefore, the proposed approach ensures a seamless integration of prior information into the score assignment, while maintaining consistency with PT when prior information is unavailable.
After obtaining the scores $s(\lambda_{i})$ for all $1\leq i\leq|\Lambda_{\text{OPT}}|$, depth assignment is done via clustering, producing disjoint subsets $\Lambda_{1},\ldots,\Lambda_{D}$. The cluster $\Lambda_{1}\subseteq\Lambda_{\text{OPT}}$ contains the hyperparameters with the highest expected reliability, and the remaining clusters $\Lambda_{2},\ldots,\Lambda_{D}$ are sorted in descending order of expected reliability. All hyperparameters in cluster $\Lambda_{d}$ are assigned depth level $d$.
Clustering can be implemented by using methods such as $K$-means or hierarchical clustering. We recommend to use agglomerative hierarchical clustering, which begins with each hyperparameter in its own cluster and iteratively merges clusters Jain & Dubes (1988).
4.2.2 Learning the Directed Edges
Having obtained the clusters $\Lambda_{1},\ldots,\Lambda_{D}$, the RG is constructed by: (i) including one node for each hyperparameter $\lambda\in\Lambda_{\text{OPT}}$; and (ii) selecting for each hyperparameter $\lambda\in\Lambda_{d}$ at depth level $d$ a subset of hyperparameters in cluster $\Lambda_{d-1}$ to serve as parents of $\lambda$ for all depth levels $2\leq d\leq K$. The resulting directed edges are intended to represent inferred reliability dependencies.
To this end, we implement feature selection via the non-negative Lasso Tibshirani (1996). Specifically, given hyperparameter $\lambda\in\Lambda_{d}$, we consider the problem of predicting the risks $\{r_{l}(Z,\lambda)\}_{l=1}^{L}$ from the risks $\{r_{l}(Z,\lambda^{\prime})\}_{l=1}^{L}$ attained by the hyperparameters $\lambda\in\Lambda_{d-1}$ at the previous depth level. The use of non-negative Lasso regression ensures that only positive correlations are represented in the DAG, preserving hierarchical reliability relationships between parent and child nodes.
Formally, using the data set $\mathcal{Z}_{\text{OPT}}$, for each hyperparameter $\lambda\in\Lambda_{d}$ we address the problem
$\min_{\beta\geq 0}$ $\sum_{Z\in\mathcal{Z}_{\text{OPT}}}\|r(Z,\lambda)-\sum_{\lambda^{\prime}\in\Lambda_{d-1}}\beta_{\lambda^{\prime}}r(Z,\lambda^{\prime})\|_{2}^{2}$ (13)
$+\tau\sum_{\lambda^{\prime}\in\Lambda_{d-1}}\beta_{\lambda^{\prime}},$
where $\beta=\{\beta_{\lambda^{\prime}}\}_{\lambda^{\prime}\in\Lambda_{d-1}}$ is the vector of non-negative regression coefficients corresponding to each potential parent node $\lambda^{\prime}\in\Lambda_{d-1}$; $r(Z,\lambda)$ is the vector containing the values $\{r_{l}(Z,\lambda)\}_{l=1}^{L_{c}}$; $\|\cdot\|_{2}$ represents the $\ell_{2}$ norm; and $\tau>0$ is a regularization parameter that controls the degree of sparsity in the solution. As for the variables $(n_{p},\{\eta_{ij}\})$ in the BT likelihood (12), the choice of the parameter $\tau$ does not affect the validity properties of RG-PT. After solving the convex problem (13), only the hyperparameters $\lambda^{\prime}\in\Lambda_{j-1}$ for which the corresponding coefficient $\beta_{\lambda^{\prime}}$ are positive are selected as parent nodes of hyperparameter $\lambda$.
4.3 FDR-Controlling Multiple Hypothesis Testing
Given the obtained RG, RG-PT performs MHT via DAGGER Ramdas et al. (2019), an FDR-controlling algorithm that operates on DAGs. Next, we first outline the reasons behind using FDR, and then explain the testing procedure.
The requirement (6) guaranteed by FWER-controlling algorithms often comes at the cost of limiting the number of hyperparameters retained in the subset $\hat{\Lambda}_{\mathcal{Z}}$. This, in turn, constrains the quality of the solutions to problem (7). Instead of constraining the probability of having any unreliable hyperparameter in set $\hat{\Lambda}_{\mathcal{Z}}$, one can control the proportion of unreliable hyperparameters in set $\hat{\Lambda}_{\mathcal{Z}}$.
Accordingly, the FDR is defined as the expected proportion of unreliable hyperparameters in set $\hat{\Lambda}_{\mathcal{Z}}$. Controlling the FDR amounts to finding a subset $\hat{\Lambda}_{\mathcal{Z}}\subseteq\Lambda$ that satisfies the inequality
$\mathbb{E}_{Z}\left[\frac{\sum_{\lambda\in\hat{\Lambda}_{\mathcal{Z}}}\mathbf{1}\{R_{l}(\lambda)>\alpha_{l}\text{ for any }l=1,\ldots,L_{c}\}}{\max(|\hat{\Lambda}_{\mathcal{Z}}|,1)}\right]\leq\delta,$ (14)
where $\mathbf{1}\{\cdot\}$ is the indicator function, and the expectation is taken over the unknown data distribution $P_{Z}$. The FDR constraint (14) contrasts with the FWER requirement (6), which aims to limit the probability of any unreliable hyperparameter in the selected set $\hat{\Lambda}_{\mathcal{Z}}$.
DAGGER is an FDR-controlling procedure that operates on a DAG. DAGGER begins testing at the root nodes of the RG, i.e., at the hyperparameters in cluster $\Lambda_{1}$, and proceeds with the clusters $\Lambda_{2},\Lambda_{3},\ldots,\Lambda_{D}$, guided by the outcomes of prior tests. If a hyperparameter is deemed unreliable, none of its descendants are tested.
Using the data set $\mathcal{Z}_{\text{MHT}}$, the $p$-values used for MHT are given by $p_{\lambda}(\mathcal{Z}_{\text{MHT}})$ in (9) for each hyperparameter $\lambda\in\Lambda_{\text{OPT}}$. The testing level $\delta_{i}$ for each hyperparameter $\lambda_{i}$, detecting $\lambda_{i}$ as reliable if $p_{\lambda_{i}}\leq\delta_{i}$, is determined based on several factors. These include the overall target FDR level $\delta$ in constraint (14), the number of reliable hyperparameters identified among those tested prior to $\lambda_{i}$, and the structure of the graph rooted at $\lambda_{i}$. We refer the reader to Appendix B and to Ramdas et al. (2019) for details.
5 Experiments
In this section, we evaluate the proposed RG-PT hyperparameter selection strategy on two problems: a sequence-to-sequence translation problem Peters & Martins (2021), and an object detection task Angelopoulos et al. (2021). Two additional experiments addressing a classical small-scale machine learning benchmark Franceschi et al. (2024) and an engineering application in the domain of telecommunications Valcarce (2020) can be found in Appendix C [^1] The codes for the experiments can be found at https://github.com/kclip/RG-PT..
Throughout, we adopt as benchmarks LTT Angelopoulos et al. (2021) and PT Laufer-Goldshtein et al. (2023). LTT is implemented by applying Benjamini-Hochberg (BH) Benjamini & Hochberg (1995) as the FDR-controlling algorithm, while PT follows Sec. 3 with the caveat that FDR-controlling FST Lynch et al. (2017) is used in lieu of an FWER-controlling scheme. LTT uses the entire calibration data set $\mathcal{Z}$ to evaluate the $p$-values used in BH, while PT and RG-PT partition $\mathcal{Z}$ into data sets $\mathcal{Z}_{\text{OPT}}$ and $\mathcal{Z}_{\text{MHT}}$. To the best of our knowledge, LTT and PT are the only existing hyperparameter selection methods that guarantee statistical validity in the sense of the FWER requirement (6) or the FDR constraint (14).
5.1 Sequence-to-Sequence Language Translation
We consider a sequence-to-sequence language translation task on the WMT16 Romanian-English dataset Bojar et al. (2016), using BLEU Papineni et al. (2002) and ROUGE-L Lin (2004) as the objectives. Following Peters & Martins (2021), the dataset is preprocessed with SentencePiece tokenization Kudo (2018), and an LSTM-based encoder-decoder is trained. Two key hyperparameters are considered:
1. The hyperparameter $\rho$ controls the sparsity of the output distribution using Entmax Peters et al. (2019), transitioning between a dense output with softmax ($\rho=1$) and sparsemax ($\rho=2$) Martins & Astudillo (2016).
2. The Fenchel-Young label smoothing strength $\epsilon$ is a training regularization hyperparameter that determines the extent to which one-hot targets are mixed with uniform noise based on Fenchel-Young losses Peters & Martins (2021). Accordingly, the original one-hot targets are assigned weight $1-\epsilon$, while the uniform distribution over all possible classes is assigned the weight $\epsilon$.
To create the initial candidate set $\Lambda$, we selected hyperparameters over a grid of 32 combinations, using 8 logarithmically spaced values in the interval $[1,2]$ for $\rho$, and values in the set $\{0.0,0.01,0.05,0.1\}$ for $\epsilon$. This selection is in line with Peters & Martins (2021).
To set up RG-PT, we leveraged the prior knowledge that less sparse settings may be more reliable than their sparser counterparts. Specifically, for any two hyperparameters $\lambda_{i}=(\rho_{i},\epsilon_{i})$ and $\lambda_{j}=(\rho_{j},\epsilon_{j})$ where $\rho_{i}<\rho_{j}$, we assigned a prior probability $\eta_{ij}=1$, reflecting this prior reliability assumption. Furthermore, the pseudocount parameter $n_{p}$, which determines the weight of prior information, was set to be equal to $|\mathcal{Z}_{\text{OPT}}|$.
Denote as $R_{\text{BLEU}}(\lambda)$ and $R_{\text{ROUGE}}(\lambda)$ the average BLEU and ROUGE-L scores, respectively, obtained for a given hyperparameter configuration $\lambda=(\rho,\epsilon)$. The goal is to guarantee the BLEU score to be above a threshold $\alpha$, while maximizing the ROUGE-L score. This amounts to an instance of problem (3), with $L=2$, $L_{c}=1$, $R_{1}(\lambda)=-R_{\text{BLEU}}(\lambda)$, $R_{2}(\lambda)=-R_{\text{ROUGE}}(\lambda)$, and $\delta=0.1$. After MHT, all the schemes choose the hyperparameter $\lambda\in\hat{\Lambda}_{\mathcal{Z}}$ with the maximum estimated value $R_{\text{ROUGE}}(\lambda)$, i.e., minimum $\hat{R}_{2}(\lambda|\mathcal{Z})$ for LTT, and minimum $\hat{R}_{2}(\lambda|\mathcal{Z}_{\text{OPT}})$ for PT and RG-PT. The data set sizes are $|\mathcal{Z}|=400$, $|\mathcal{Z}_{\text{OPT}}|=200$, and $|\mathcal{Z}_{\text{MHT}}|=200$.
Fig. 5 illustrates the ROUGE-L score achieved on the test data by each calibration method, plotted against the target value for the BLEU score. The results demonstrate that RG-PT consistently maintains higher ROUGE-L scores, even under stricter requirements for the BLEU score. This highlights RG-PT’s advantage in effectively exploring the hyperparameter space, enabling a more efficient testing procedure and identifying superior hyperparameter configurations that still statistically satisfy the desired conditions on the risk functions.
Figure 5: Test ROUGE-L scores achieved by LTT, PT, and RG-PT methods as a function of the target reliability value for the BLEU score.
5.2 Image Segmentation for Object Detection
We now evaluate the proposed RG-PT framework on a multi-objective image segmentation task for object detection, leveraging the MS-COCO dataset Lin et al. (2014) and a pretrained detector from Detectron2 Wu et al. (2019) as done in Angelopoulos et al. (2021). The task involves three distinct objectives: (i) detecting objects within an image (object detection); (ii) delineating object boundaries (image segmentation); and (iii) assigning correct labels to detected objects (object classification). These tasks are measured using recall, intersection-over-union (IoU), and classification accuracy, respectively. The goal is to control classification errors while optimizing recall and segmentation quality, addressing the trade-offs among these objectives.
The performance of the detection is determined by three hyperparameters:
1. The object recall threshold ($\lambda_{1}$) controls the threshold for selecting objects based on confidence scores. Reducing the value of $\lambda_{1}$ lowers the confidence threshold, which allows more objects to be selected at the cost of, potentially, increasing false positives.
2. The mask size threshold ($\lambda_{2}$) tunes the size of the bounding masks used to segment objects, impacting the IoU score.
3. The classification certainty level ($\lambda_{3}$) controls the certainty level required for object classification, adjusting the tolerance for inclusion in the set of labels assigned to each detected object.
Denote as $R_{1}(\lambda)$, $R_{2}(\lambda)$, and $R_{3}(\lambda)$ the risks associated with recall, IOU, and coverage, respectively, for hyperparameter $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3})$. Controlling these risks in the context of problem (3) is equivalent to having $L=3$ and $L_{c}=3$. Additionally, as in Angelopoulos et al. (2021), we set the targets as $\alpha_{1}=0.5$, $\alpha_{2}=0.5$, and $\alpha_{3}=0.75$ with $\delta=0.1$. Within the set $\hat{\Lambda}_{\mathcal{Z}}$ of reliable hyperparameters returned by the algorithm of choice, we choose the hyperparameter in subset $\hat{\Lambda}_{\mathcal{Z}}$ with the lowest value of $\lambda_{1}$ in order to increase the number of detected objects as much as possible. No prior knowledge was leveraged in creating the RG in this experiment by setting $n_{p}=0$.
We compare the distribution of the hyperparameters returned by LTT, PT, and RG-PT. The distribution is obtained by running 200 trials for each algorithm over different splits of calibration data $\mathcal{Z}$ into subsets $\mathcal{Z}_{\text{OPT}}$ and $\mathcal{Z}_{\text{MHT}}$ with $|\mathcal{Z}_{\text{OPT}}|=1500$ and $|\mathcal{Z}_{\text{MHT}}|=1500$. As shown in Fig. 6, the results demonstrate that RG-PT tends to return lower values for $\lambda_{1}$ than both LTT and PT. In particular, both the mean and dispersion for RG-PT are lower than those for LTT and PT. A lower threshold $\lambda_{1}$ allows the detector to select more objects, which directly enhances object recall, while still maintaining controlled levels of segmentation and classification accuracy (see also Fig. 2).
Figure 6: Example of of hyperparameter distributions for LTT, PT, and RG-PT methods. The box plots show the range (denoted by the horizontal black whiskers), median (represented by the thick black vertical lines), and interquartile range (depicted by the boxes) of the object recall threshold hyperparameter $\lambda_{1}$.
6 Conclusion
In this paper, we have introduced RG-PT, a novel framework for multi-objective hyperparameter selection that integrates MHT with the concept of RGs to capture interdependencies among candidate hyperparameters. By leveraging a DAG structure informed by prior knowledge and data, RG-PT enables a more powerful parallel testing of hyperparameters compared to the state-of-the-art methods LTT and PT. RG-PT provides statistical guarantees through FDR control, while expanding the space of reliable hyperparameter configurations, leading to a superior optimization of auxiliary objectives.
Future work may investigate the use of synthetic data for the derivation of an RG, as well as the integration with sequential testing methods based on e-processes Zecchin & Simeone (2024).
Acknowledgements
This work was supported by the European Union’s Horizon Europe project CENTRIC (101096379), by the Open Fellowships of the EPSRC (EP/W024101/1), and by the EPSRC project (EP/X011852/1).
Appendix A Fixed Sequence Testing
In this section, we provide a brief overview of FST for controlling the FDR, which we used in our simulations for PT. While PT, as outlined in Laufer-Goldshtein et al. (2023), is designed to support control of the FWER, our focus in this paper is on FDR control.
MHT methods such as the Bonferroni correction for FWER control Rice (2006) and the Benjamini-Yekutieli (BY) procedure for FDR control Benjamini & Yekutieli (2001) do not leverage any side information about the relative reliability of the hyperparameters. When such information is available during calibration, FST can be used to test hyperparameters in order of expected reliability. When the ordering information is accurate, FST can be beneficial to reduce the FDR Lynch et al. (2017). In this section, we briefly describe the FST procedure for FDR control.
With FST, the candidate hyperparameters are ordered as $\lambda_{(1)},\ldots,\lambda_{(|\Lambda|)}$ using side information. The ordering ideally lists the hyperparameters from the most to the least likely to meet the reliability criterion (2).
Starting with $i=1$, each hyperparameter $\lambda_{(i)}$ is tested sequentially based on its $p$-value $p_{\lambda_{(i)}}$ against an adjusted critical value $\delta_{i}$ that decreases with the index $i=1,2,\ldots,|\Lambda|$. At each step $i$, the hyperparameter $\lambda_{(i)}$ is deemed to be reliable if $p_{\lambda_{(i)}}\leq\delta_{i}$. Testing continues until $k$ hyperparameters are deemed to be unreliable, at which point testing stops. The choice of the integer $k$ is typically set as a small proportion, often around 5-10%, of the total number of hypotheses, $|\Lambda|$.
The critical values $\delta_{i}$ are adapted to account for the position of each hypothesis in the testing sequence. These values are specifically designed to control the FDR under various dependency structures among the $p$-values. For the case of interest here, which is arbitrary dependence of the $p$-values, the critical levels can be set as Lynch et al. (2017)
$\delta_{i}=\begin{cases}\frac{\delta}{k}&\quad\text{if}\quad i\leq k\\ \frac{(|\Lambda|-k+1)\delta}{(|\Lambda|-i+1)k}&\quad\text{if}\quad i>k,\end{cases}$ (15)
where $\delta$ is the target FDR level and $k$ is the number of unreliable hyperparameters allowed before testing stops.
The final set of reliable hyperparameters is $\hat{\Lambda}_{\mathcal{Z}}=\{\lambda_{(1)},\ldots,\lambda_{(j)}\}$, where $j$ corresponds to the index of the last hyperparameter tested before stopping. This ensures that the FDR is rigorously maintained below level $\delta$.
Appendix B Summary of DAGGER
This section outlines the step-up procedure used in DAGGER Ramdas et al. (2019) to dynamically adjust the testing thresholds. DAGGER determines the testing thresholds adaptively based on the structure of the DAG and on the outcomes of previously tested hypotheses. At each depth level of the DAG, thresholds are updated dynamically to control the FDR, while respecting the hierarchical dependencies encoded by the DAG.
At each depth $d$, only the hyperparameters with no unreliable parents are considered for testing. If any parent of a hyperparameter $\lambda$ is deemed unreliable by DAGGER, all of its descendants are also automatically deemed unreliable. The threshold for testing the $i$-th hypothesis at depth $d$ is given by
$\delta_{i}(r)=\frac{v_{i}}{V}\cdot\frac{\delta}{\beta(m_{i}+r+R_{1:d-1}-1)},$ (16)
where $r\in[1,|\Lambda_{d}|]$ is a parameter set as detailed below; $v_{i}$ is the effective number of leaves in the subgraph rooted at the current node; $V$ is the total number of leaves in the DAG; $m_{i}$ is the effective number of nodes in the subgraph rooted at the current node; $R_{1:d-1}$ is the total number of rejections at depths $1$ through $d-1$; and $\beta(\cdot)$ is a reshaping function, such as the Benjamini-Yekutieli Benjamini & Yekutieli (2001) function $\beta_{BY}(x)=x\sum_{k=1}^{V}1/k$, which is designed to ensure FDR control under arbitrary dependence.
The effective number of leaves $v_{i}$ and the effective number of nodes $m_{i}$ for node $i$ are calculated as follows. If $i$ is a leaf, then we have $v_{i}=m_{i}=1$. Otherwise, the values $v_{i}$ and $m_{i}$ are calculated recursively from leaves to roots as
$v_{i}=\sum_{j\in\text{children}(i)}\frac{v_{j}}{|\text{parents}(j)|},$ (17)
and
$m_{i}=1+\sum_{j\in\text{children}(i)}\frac{m_{j}}{|\text{parents}(j)|},$ (18)
where $\text{children}(i)$ and $\text{parents}(i)$ denote the sets of the children and the parents of node $i$, respectively.
The parameter $r$ at depth $d$ needs to be determined before testing can begin. To maximize the number of rejections while ensuring FDR control, DAGGER calculates the value
$R=\arg\max_{r=1,\ldots,|\Lambda_{d}|}\left[\sum_{\lambda_{i}\in\Lambda_{d}}\mathbf{1}\{p_{\lambda_{i}}(\mathcal{Z}_{\text{MHT}})\leq\delta_{i}(r)\}\geq r\right].$ (19)
The threshold $\delta_{i}(R)$ is then used to perform the testing for hyperparameter $\lambda_{i}$.
The step-up procedure (16) ensures that thresholds are increasingly relaxed, i.e., increased, as we move further down the DAG. The overall algorithm is described in Algorithm 1, and an example illustrating DAGGER’s operation is shown in Fig. 7. The figure highlights how thresholds are updated and decisions propagated through the DAG. For simplicity, the figure assumes the reshaping function $\beta(\cdot)$ to be the identity function $\beta(x)=x$. First, the values for the effective number of leaves $v_{i}$ and the effective number of nodes $m_{i}$ are calculated for each hyperparameter $\lambda_{i}$. Next, going level by level, the thresholds $\delta_{i}(R)$ are calculated using (16) and (19), and each hyperparameter $\lambda_{i}$ is tested by comparing $p_{\lambda_{i}}(\mathcal{Z}_{\text{MHT}})$ with $\delta_{i}(R)$.
Figure 7: Illustration of the DAGGER algorithm’s operation to control the FDR at $\delta=0.1$. At each step, a hyperparameter is tested, starting from the root nodes and progressing level by level through the DAG. The testing thresholds $\delta_{i}$ are computed for each hyperparameter $\lambda_{i}$ using the step-up procedure in (16) and (19), using the identity function as $\beta(\cdot)$. The $p$-value of each hyperparameter is compared against its respective threshold $\delta_{i}$ to assess the reliability of $\lambda_{i}$.
Algorithm 1 DAGGER Ramdas et al. (2019)
  Input: DAG structure, $p$-values $\{p_{\lambda}(\mathcal{Z}_{\text{MHT}})\}$, target FDR level $\delta$
  Output: Set of reliable hyperparameters $\hat{\Lambda}_{\mathcal{Z}}$
  for depth $d=1,\dots,D$ do
     for each hyperparameter $\lambda_{i}$ in cluster $\Lambda_{d}$ do
        if all parent hyperparameters of $\lambda_{i}$ are deemed as reliable then
           Evaluate threshold $\delta_{i}(R)$ using (16) and (19)
           if $p_{\lambda_{i}}(\mathcal{Z}_{\text{MHT}})\leq\delta_{i}(R)$ then
              Detect $\lambda_{i}$ as reliable
           else
              Detect $\lambda_{i}$ as unreliable
           end if
        end if
     end for
     Update $\hat{\Lambda}_{\mathcal{Z}}$ with all the hyperparameters detected as reliable at depth $d$
  end for
  return  $\hat{\Lambda}_{\mathcal{Z}}$
Appendix C Additional Experiments
C.1 Image Classification
In this experiment, following Franceschi et al. (2024), we consider the problem of hyperparameter selection for a support vector machine (SVM) model used to classify images from the Fashion MNIST dataset. The Fashion MNIST is a widely used benchmark for image classification, consisting of 70,000 grayscale images of 10 different clothing categories.
We consider two risk functions ($L=2$) in problem (3), namely the classification error $R_{\text{err}}(\lambda)$ and the recall, $R_{\text{rec}}(\lambda)$. The classification error $R_{\text{err}}(\lambda)$ measures the proportion of incorrectly classified images out of the total number of samples. The recall measures the ability of a model to correctly identify all the relevant instances of each class. Accordingly, for each class, the recall is computed as the ratio of correctly identified instances of that class to the total number of actual instances of the same class in the dataset. The recall $R_{\text{rec}}(\lambda)$ represents the average of the recall values across all classes.
With reference to problem (3), we aim at minimizing recall, i.e. $R_{2}(\lambda)=R_{\text{rec}}(\lambda)$, while keeping the classification error rate below $0.3$, i.e. $R_{1}(\lambda)=R_{\text{acc}}(\lambda)$, $L_{c}=1$, and $\alpha_{1}=0.3$. The goal is therefore defined as
$\min_{\lambda\in\Lambda}R_{\text{rec}}(\lambda)\quad\text{subject to}\quad R_{\text{acc}}(\lambda)<0.3.$ (20)
This is a non-trivial problem since the accuracy maximizing model may not also optimize the recall Powers (2011).
The SVM model requires the selection of two hyperparameters Franceschi et al. (2024). The regularization parameter, $C$, controls the desired trade-off between maximizing the margin and minimizing the classification error. Lower values of $C$ allow for a softer margin that can overlook some misclassification errors, while higher values enforce stricter classification error requirements. The kernel coefficient, $\gamma$, determines the impact of a single training example on the decision boundary, with higher values capturing finer details in the data set but risking overfitting. To create the initial candidate set $\Lambda$, we selected hyperparameters over a grid of 25 combinations, using five logarithmically spaced values in the intervals $[-3,3]$ and $[-4,1]$ for $C$ and $\gamma$, respectively. This selection is in line with approaches such as Yogatama et al. (2015) for SVM hyperparameter selection.
We used 5,000 data points for training the SVM, and used an additional 5,000 data points as calibration data $\mathcal{Z}$. The calibration data set $\mathcal{Z}$ was in turn divided into two groups of size 2,500, for the data sets $\mathcal{Z}_{\text{OPT}}$ and $\mathcal{Z}_{\text{MHT}}$, respectively.
Fig. 8 illustrates the testing procedures of LTT Angelopoulos et al. (2021), PT Laufer-Goldshtein et al. (2023), and RG-PT. The x- and y-axes represent the logarithmic scales of the two hyperparameters $C$ and $\gamma$, while the contours indicate levels of recall $R_{\text{rec}}(\lambda)$, evaluated on the test data set, as a function of the hyperparameters $\lambda=(c,\gamma)$. The numbers illustrate the testing order for each testing method. Note that than LTT, which uses BY, does not follow any inferred order on the hyperparameters, and thus does not have the order labels in the figures. Furthermore, while LTT and PT test hyperparameters one by one, following a linear trajectory, RG-PT proceeds along a DAG, testing at the same time all hyperparameters at the same depth in the DAG.
LTT and PT are seen to stop at the sixth tested hyperparameter, yielding the set of reliable hyperparameters $\hat{\Lambda}_{\mathcal{Z}}$ marked as green dots. In contrast, RG-PT returns a much larger set $\hat{\Lambda}_{\mathcal{Z}}$ of reliable hyperparameters, also marked as green dots. Choosing within these sets the hyperparameter that minimizes the estimated recall as per problem (20) yields the solutions indicated as green stars, corresponding to a test recall of 0.727 for LTT and PT, and 0.332 for RG-PT.
It is important to note that all three methods yield test accuracies below the 0.3 threshold in (20). Specifically, the hyperparameters selected by LTT and PT result in a test accuracy error of 0.267, while those chosen by RG-PT achieve a slightly higher accuracy error of 0.286. Although the accuracy of RG-PT is closer to the threshold, it remains consistent with the statistical guarantee outlined in (20). In fact, RG-PT achieves a lower recall while maintaining the desired accuracy constraint, whereas LTT and PT follow a more conservative approach, leading to a reliable hyperparameter with higher recall.
(a) LTT
(b) PT
(c) RG-PT
Figure 8: Illustration of the hyperparameter selection procedure followed by LTT (a), PT (b), and RG-PT (c) for the setting studied in Sec. C.1. Each node represents a hyperparameter $\lambda=(C,\gamma)$, with the numbers representing the testing order. Green nodes show the hyperparameters included in the reliable set $\hat{\Lambda}_{\mathcal{Z}}$, and the star node shows the hyperparameter in set $\hat{\Lambda}_{\mathcal{Z}}$ with the lowest recall rate.
C.2 Radio Access Scheduling
In this section, we study a telecommunications engineering problem, namely the optimization of a radio access scheduler Valcarce (2020). In this setup, each user equipment (UE) belongs to one of four quality-of-service (QoS) classes, assigned at random, each with its own delay and bit rate requirements de Sant Ana & Marchenko (2020). The goal is to control the delay of UEs in a given QoS class, while simultaneously minimizing the delays for UEs in the other three QoS classes.
Accordingly, in the context of problem (3), we choose $L=4$ and $L_{c}=1$, and we set the risk $R_{i}(\lambda)$ to be equal to the average delay of QoS class $i$ for $1\leq i\leq 4$. We aim to keep $R_{2}(\lambda)$ below 15 ms, while minimizing $R_{1}(\lambda)$, $R_{3}(\lambda)$, and $R_{4}(\lambda)$. Formally, the problem is stated as
$\min_{\lambda\in\Lambda}\{R_{1}(\lambda),R_{3}(\lambda),R_{4}(\lambda)\}\;\text{subject to}\;R_{2}(\lambda)<15\;\text{ms}.$ (21)
The scheduling algorithm at the base station allocates spectral resources to the UEs. As in Valcarce (2020), the UEs are randomly distributed within a $1\,\text{km}^{2}$ area containing a centrally located base station. Each UE has an initial buffer of 100 packets, and moves at random speeds and directions. Resource allocation is carried out in intervals of 1 ms, called transmission time intervals (TTIs), over 10,000 TTIs per episode.
The resource allocation algorithm is controlled by a set of hyperparameters $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})\in\Lambda$, where each $\lambda_{i}$ adjusts a specific criterion in the reward model as detailed in Valcarce (2020). Hyperparameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ determine respectively the channel quality for each UE, the total queue sizes at the UEs, the age of the oldest packet in each UE’s buffer, and the fairness in resource block allocation among UEs.
Calibration and test data were generated using the Nokia wireless suite Valcarce (2020). For each run, we used 100 episodes for calibration and 100 episodes for testing.
The candidate hyperparameter set $\Lambda$ was generated by keeping $\lambda_{1}$ and $\lambda_{2}$ at the values $\lambda_{1}^{*}$ and $\lambda_{2}^{*}$ recommended by de Sant Ana & Marchenko (2020), and linearly sweeping hyperparameters $\lambda_{3}$ and $\lambda_{4}$ in [0.02, 0.2] and [-0.1, 0.1], respectively, with 10 steps each, resulting in a total of 100 combinations.
Fig. 9 presents the results of using PT and RG-PT to optimize the hyperparameter $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$. Both methods successfully meet the statistical guarantee of $R_{2}(\lambda)<15$ ms for class 2. However, RG-PT demonstrates a greater ability to explore the hyperparameter space $\Lambda$, identifying configurations that more effectively minimize the average delay across the other three classes.
Figure 9: Distribution of the average delay for the four QoS classes using hyperparameters optimized by PT (left column) and RG-PT (right column). The dashed red line indicates the target threshold for the average delay in QoS class 2.